Due to a recession, expected inflation this year is only 1.0%. However, the inflation rate in Year 2 and thereafter is expected to be constant at some level above 1.0%. Assume that the expectations theory holds and the real risk-free rate (r*) is 2.0%. If the yield on 10-year Treasury bonds equals the 1-year yield plus 15.00%, what inflation rate is expected after Year 1?

A) 17.667

B) 16.000

C)18.000

D) 15.00
Answer #1

Yield on 1 year treasury securities (rt1)= r* + IP1
= 2% + 1%
= 3%

Yield on 10 year treasury securities (rt10)= rt1+15%
= 3% + 15%
= 18%

rt10=r*+IP
Thus, IP = 18% - 2% = 16%

Average inflation premium (IP)= 16%

16% = (IP1+IP2+IP3+IP4+IP5+IP6+IP7+IP8+IP9+IP10) / 10

16 * 10 = 1% + X+X+X+X+X+X+X+X+X

160 - 1 = 9X

159 = 9X

X = 159 / 9

X = 17.667%

Inflation rate is expected after year 1= 17.667%